Mathematics 242 – Principles of Analysis Definitions Quiz 1 – September 9, 2005

Your Name:

In both questions here, x represents a real number.

1) Give the converse of the implication: “If |x − 2| ≤ 3, then −1 ≤ x ≤ 1.” Is the converse a true statement?

2) Give the contrapositive form of the implication from 1. Is the contrapositive a true statement?